# Generalized Categories Discovery for Long-tailed Recognition

### v1

@@ 4.1 Setup
Table 2: Performance comparison on CIFAR100 and ImageNet100 with varying imbalance factors $\rho$ ($\rho=0.5$ for original GCD setting). Un1 represents unknown-aware accuracies, while Un2 refers to unknown-agnostic accuracies. Our method consistently excels in novel class performance under imbalanced settings and remains competitive in balanced settings. Mean and standard deviation of clustering accuracy are reported over 3 runs for multiple methods. Higher means are bolded, while results within one standard deviation are not.
IMF Metrics Methods
CIFAR100 ImageNet100
GCD ORCA OpenCon ImbaGCD Ours GCD ORCA OpenCon ImbaGCD Ours
$\rho=0.5$ All 45.41$\pm$0.13 55.68$\pm$0.33 51.85$\pm$0.63 53.51$\pm$0.26 52.20$\pm$0.30 77.12$\pm$0.56 74.93$\pm$0.34 82.22$\pm$0.56 81.90$\pm$0.67 81.06$\pm$0.39
Known 67.61$\pm$0.12 66.41$\pm$0.31 69.07$\pm$0.29 68.09$\pm$0.13 70.45$\pm$0.31 87.02$\pm$0.36 89.21$\pm$0.06 90.65$\pm$0.04 91.19$\pm$0.15 91.03$\pm$0.03
Un1 34.31$\pm$0.22 42.63$\pm$0.67 45.76$\pm$0.32 47.92$\pm$0.33 46.96$\pm$0.35 57.48$\pm$0.76 67.18$\pm$0.27 78.12$\pm$0.80 77.89$\pm$0.88 77.22$\pm$0.25
Un2 18.12$\pm$0.34 38.95$\pm$0.76 42.11$\pm$0.54 46.22$\pm$0.33 46.25$\pm$0.25 42.68$\pm$0.77 65.43$\pm$0.37 78.01$\pm$0.83 77.79$\pm$0.97 77.68$\pm$0.30
$\rho=1$ All 48.36$\pm$0.08 47.37$\pm$0.52 53.20$\pm$0.33 54.06$\pm$0.45 55.82$\pm$0.40 63.06$\pm$0.66 68.01$\pm$0.31 82.44$\pm$0.24 82.34$\pm$0.39 82.28$\pm$0.68
Known 71.48$\pm$0.24 65.14$\pm$0.17 68.00$\pm$0.06 67.98$\pm$0.37 68.00$\pm$0.35 88.80$\pm$0.43 88.99$\pm$0.05 90.62$\pm$0.12 90.56$\pm$0.18 91.30$\pm$0.10
Un1 25.24$\pm$0.07 34.93$\pm$1.04 43.58$\pm$0.34 43.39$\pm$0.59 44.54$\pm$0.40 37.29$\pm$1.11 47.68$\pm$0.70 74.45$\pm$0.36 74.56$\pm$0.79 75.06$\pm$1.00
Un2 12.02$\pm$0.46 29.61$\pm$1.02 38.40$\pm$0.65 38.76$\pm$0.9 42.65$\pm$0.89 31.16$\pm$1.48 47.28$\pm$0.66 74.35$\pm$0.38 74.24$\pm$0.89 73.39$\pm$1.38
$\rho=5$ All 63.13$\pm$0.12 56.37$\pm$0.12 61.84$\pm$0.29 59.48$\pm$0.48 62.85$\pm$0.53 77.29$\pm$0.17 76.79$\pm$0.13 81.87$\pm$0.23 83.01$\pm$0.11 86.14$\pm$0.08
Known 70.96$\pm$0.13 64.40$\pm$0.11 69.37$\pm$0.23 67.82$\pm$0.06 70.15$\pm$0.05 87.51$\pm$0.21 89.02$\pm$0.12 90.77$\pm$0.10 88.89$\pm$0.06 91.69$\pm$0.31
Un1 24.04$\pm$0.12 25.36$\pm$0.40 35.06$\pm$0.34 37.87$\pm$1.59 36.72$\pm$1.23 23.83$\pm$0.24 20.19$\pm$0.55 39.76$\pm$1.4 54.35$\pm$0.38 61.17$\pm$0.91
Un2 7.18$\pm$0.34 16.18$\pm$0.23 24.21$\pm$0.67 27.64$\pm$2.26 26.36$\pm$1.98 14.80$\pm$0.23 16.42$\pm$0.73 37.93$\pm$1.3 53.97$\pm$0.38 59.10$\pm$0.59
$\rho=10$ All 66.21$\pm$0.23 59.61$\pm$0.27 65.05$\pm$0.29 63.21$\pm$0.05 62.84$\pm$0.06 83.32$\pm$0.15 81.98$\pm$0.03 84.60$\pm$0.10 83.23$\pm$0.17 87.44$\pm$0.31
Known 70.36$\pm$0.31 64.27$\pm$0.32 69.80$\pm$0.23 67.82$\pm$0.07 66.78$\pm$0.04 89.71$\pm$0.03 89.24$\pm$0.02 90.85$\pm$0.12 87.48$\pm$0.23 91.51$\pm$0.11
Un1 24.74$\pm$0.45 26.21$\pm$0.69 33.01$\pm$0.34 34.87$\pm$0.70 34.08$\pm$0.54 22.17$\pm$0.18 16.94$\pm$0.07 26.70$\pm$0.11 42.62$\pm$0.50 51.18$\pm$2.53
Un2 7.28$\pm$0.21 13.01$\pm$0.34 17.57$\pm$0.67 21.68$\pm$0.29 23.52$\pm$0.32 11.47$\pm$0.28 10.31$\pm$0.22 22.89$\pm$0.27 41.23$\pm$0.58 47.61$\pm$2.59
@@ 4.2 Compared with SOTA
We benchmark our method against four leading GCD techniques: (1) GCD [18] - a foundational GCD framework that clusters unknown categories without exhaustive labeling. (2) ORCA [1] - melds supervised and unsupervised learning, addressing open-world class imbalances. (3) OpenCon [17] - uses contrastive learning to adapt to new categories with minimal supervision. (4) ImbaGCD [11] - employs the EM algorithm, combining Sinkhorn-based pseudo-labeling with contrastive updates. The class prior and prototype are refreshed every epoch. Our evaluation spans original GCD, balanced, and various imbalanced scenarios for a thorough comparison.
@@
Figure 1: Performance trends of Overall, Known, Unknown-Aware, and Unknown-Agnostic accuracies across different $\beta$ values with $\alpha=0$ constant on CIFAR100 with $rho=5$.

### v2

@@ 4.1 Setup
Table 2: Performance comparison on CIFAR100 with varying imbalance factors $\rho$ ($\rho=0.5$ for original GCD setting). Un1 represents unknown-aware accuracies, while Un2 refers to unknown-agnostic accuracies. Our method consistently excels in novel class performance under imbalanced settings and remains competitive in balanced settings. Mean and standard deviation of clustering accuracy are reported over 3 runs for multiple methods. Higher means are bolded, while results within one standard deviation are not.
IMF Metrics CIFAR100
GCD ORCA OpenCon ImbaGCD Ours
$\rho=0.5$ All 45.41$\pm$0.13 55.68$\pm$0.33 51.85$\pm$0.63 53.51$\pm$0.26 52.20$\pm$0.30
Known 67.61$\pm$0.12 66.41$\pm$0.31 69.07$\pm$0.29 68.09$\pm$0.13 70.45$\pm$0.31
Un1 34.31$\pm$0.22 42.63$\pm$0.67 45.76$\pm$0.32 47.92$\pm$0.33 46.96$\pm$0.35
Un2 18.12$\pm$0.34 38.95$\pm$0.76 42.11$\pm$0.54 46.22$\pm$0.33 46.25$\pm$0.25
$\rho=1$ All 48.36$\pm$0.08 47.37$\pm$0.52 53.20$\pm$0.33 54.06$\pm$0.45 55.82$\pm$0.40
Known 71.48$\pm$0.24 65.14$\pm$0.17 68.00$\pm$0.06 67.98$\pm$0.37 68.00$\pm$0.35
Un1 25.24$\pm$0.07 34.93$\pm$1.04 43.58$\pm$0.34 43.39$\pm$0.59 44.54$\pm$0.40
Un2 12.02$\pm$0.46 29.61$\pm$1.02 38.40$\pm$0.65 38.76$\pm$0.9 42.65$\pm$0.89
$\rho=5$ All 63.13$\pm$0.12 56.37$\pm$0.12 61.84$\pm$0.29 59.48$\pm$0.48 62.85$\pm$0.53
Known 70.96$\pm$0.13 64.40$\pm$0.11 69.37$\pm$0.23 67.82$\pm$0.06 70.15$\pm$0.05
Un1 24.04$\pm$0.12 25.36$\pm$0.40 35.06$\pm$0.34 37.87$\pm$1.59 36.72$\pm$1.23
Un2 7.18$\pm$0.34 16.18$\pm$0.23 24.21$\pm$0.67 27.64$\pm$2.26 26.36$\pm$1.98
$\rho=10$ All 66.21$\pm$0.23 59.61$\pm$0.27 65.05$\pm$0.29 63.21$\pm$0.05 62.84$\pm$0.06
Known 70.36$\pm$0.31 64.27$\pm$0.32 69.80$\pm$0.23 67.82$\pm$0.07 66.78$\pm$0.04
Un1 24.74$\pm$0.45 26.21$\pm$0.69 33.01$\pm$0.34 34.87$\pm$0.70 34.08$\pm$0.54
Un2 7.28$\pm$0.21 13.01$\pm$0.34 17.57$\pm$0.67 21.68$\pm$0.29 23.52$\pm$0.32
@@ 4.2 Compared with SOTA
We benchmark our method against four leading GCD techniques: (1) GCD [18], a foundational GCD framework that clusters unknown categories without exhaustive labeling. (2) ORCA [1], melds supervised and unsupervised learning, addressing open-world class imbalances. (3) OpenCon [17], uses contrastive learning to adapt to new categories with minimal supervision. (4) ImbaGCD [11], employs the EM algorithm, combining Sinkhorn-based pseudo-labeling with contrastive updates. The class prior and prototype are refreshed every epoch. Our evaluation spans original GCD, balanced, and various imbalanced scenarios for a thorough comparison.
@@
Figure 1: Performance trends of Overall, Known, Unknown-Aware, and Unknown-Agnostic accuracies across different $\beta$ values with $\alpha=0$ constant on CIFAR100 with $\rho=5$.